# ByteCheckpoint: A Unified Checkpointing System for Large Foundation Model Development

### v2

@@ 3.1 Unified Architecture
We present ByteCheckpoint, a unified checkpointing system for LFM development that supports multiple training frameworks and storage backends. Fig. 5 illustrates the system architecture.
API. ByteCheckpoint’s APIs provide a unified interface for user code across various training frameworks. To save checkpoints, users first prepare the corresponding training states, checkpoint path, framework name, and performance-related options, then call bytecheckpoint.save on each training worker. This high-level interface abstracts underlying system complexities, such as device mesh configurations, save/reshard plan generation, and I/O operations. The API is compatible with numerous frameworks and storage backends, and remains independent of the parallelism setting.
Planner. The Planner receives arguments (training states, checkpoint path, etc.) from the API layer and determines the tensors and other states to save or load for each training worker, based on the parallelism configuration and the worker’s rank. Each worker leverages the Planner to initially create local saving and loading plans, and subsequently collaborates to create global plans. Given the framework-specific specifications, such as distributed tensor abstractions and device mesh configurations, we implement a tailored planner for each training framework to extract information from these specifications and generate plans.
@@
Storage I/O. Analogous to the design of the Planner layer, the Storage I/O layer encapsulates different storage backends and manages backend-specific read/write operations and optimizations. The interface between the Engine layer and the Storage I/O layer remains unified across different storage backends, facilitating the seamless integration of new storage backends. ByteCheckpoint supports multiple storage options, including in-memory checkpoint storage [52], local disk storage, and remote storage systems (e.g., HDFS and NAS).
@@
Fig. 6 illustrates a typical use case of ByteCheckpoint. Initially, users define a dictionary specifying the states to be saved or loaded, encompassing the model, optimizer, dataloader, and additional states. Checkpoint operations are subsequently performed using bytecheckpoint.save and bytecheckpoint.load API calls. Specifically, at the outset of training resumption, stage transition, or evaluation, bytecheckpoint.load is invoked to load saved distributed checkpoints. Checkpoint resharding occurs automatically during loading when there are changes in parallelism. Throughout training, bytecheckpoint.save is periodically invoked to save checkpoints.
@@ 3.2 Decoupled Checkpoint Representation
Tensor representation. Each tensor is uniquely identified by a "fully qualified name" (FQN) and has a global shape, representing its original shape before sharding. Tensors can be either sharded or replicated across ranks (also known as training workers). For sharded tensors, the specific shard held by a worker is determined by three factors: the parallelism (e.g., tensor parallelism) applied to the tensor, the sharding dimension (e.g., dimension 1 for an MLP weight tensor), and the group rank within the corresponding parallelism group. Based on a tensor’s FQN, global shape, and sharding specification, ByteCheckpoint creates the tensor shard metadata (ShardMeta) for each rank, which is utilized in checkpoint storage to represent the tensor shard, independent of parallelism settings. More precisely, ShardMeta for a tensor shard is an index tuple (fqn, nD_offsets, nD_lengths), where nD_offsets and nD_lengths indicate the local shard’s offsets and lengths along the original multi-dimensional axes in its global shape.
@@
Checkpoint Representation. Fig. 7 illustrates the checkpoint representation in ByteCheckpoint. Distributed checkpoints comprise a global metadata file and multiple storage files. Each rank invokes ByteCheckpoint to generate three distinct files: a model state file, an optimizer state file, and an extra state storage file. The Dataloader state file is generated only by training workers whose ranks for all parallelism degrees, except for DP degrees, are 0. For tensor shards in model and optimizer states, their Metadata primarily consists of three parts: BasicMeta, which records essential information of individual tensor shards such as stride and device, critical for recovering tensor runtime states; ShardMeta, as previously introduced, which records the relative position information of tensor shards in the complete tensor; ByteMeta, which specifies the byte start offset and length of each tensor shard within the storage file. All tensor metadata are consolidated into the global metadata file, and a mapping, termed TensorShardtoBasicByteMap, is established between saved tensor shards and storage files based on the ShardMeta, BasicMeta, and ByteMeta of each tensor shard, ensuring accurate data retrieval. The global metadata file also includes a LoaderShardtoByteMap, which records the file index information of sharded states in each dataloader. All storage files and the global metadata file are stored in the specified storage backend designated by the checkpoint path when bytecheckpoint.save is invoked.
Figure 7: Checkpoint representation in ByteCheckpoint. The pipeline parallelism is applied in this example. Dataloader states are saved by rank 0.
@@ 3.3 Checkpoint Resharding Workflow
Figure 8: An illustration of online resharding workflow. Checkpoint $k$ denotes the distributed checkpoint saved by a previous active worker $k$.
ByteCheckpoint implements a generic workflow for saving and loading, automatically resharding checkpoints during the loading process. Leveraging the independence between the Engine Layer, Planner Layer, and Storage I/O Layer, we can execute consistent saving and loading steps across different training frameworks and storage backends. We illustrate the checkpoint resharding workflow below and in Fig. 8, emphasizing how our checkpoint representation enables flexible load-time resharding for GPU states. The process for checkpoint saving and loading without resharding follows a similar procedure.
Step 1. To initiate checkpoint loading with resharding, each rank invokes the bytecheckpoint.load() API, specifying the checkpoint path along with the model, optimizer, dataloader, and extra states to be restored. All ranks then load the global metadata file from the designated checkpoint path.
Step 2. For each tensor shard in the given model and optimizer, each rank queries the TensorShardToBasicByteMap within the global metadata file, identifying matching segments between the saved tensor shards and the sharding specification of new shards. After identifying these matches, the planner constructs a local loading plan, which includes BasicMeta and ByteMeta tailored to the specific shard. This identification mechanism is illustrated at the bottom of Fig. 8.
@@
Step 5. The execution engine on each rank selects the storage backend wrapper in the Storage I/O layer based on the checkpoint path then executes the loading pipeline (Sec. 4.2).
Step 6. Upon completion of loading, each rank utilizes the optimized asynchronous collective barrier primitive ensuring the atomicity of the distributed loading (Sec. 5.2).
Figure 9: Examples of dataloader resharding. Sharded states such as token buffers are resharded according to the changes of parallelism configurations.
Dataloader resharding. Similar to tensor resharding in Fig. 8, ByteCheckpoint reshards the dataloader by querying the LoaderShardtoByteMap. The illustration of dataloader resharding is depicted in Fig. 9. When parallelism configuration changes, for dataloader checkpoints, replicated states can be loaded directly, whereas sharded states such as token buffers require resharding. Specifically, when the DP degree size remains constant while other parallel degrees are altered (TP degree changes in the example of Fig. 9), the token buffers should be copied to the destination workers. When there is a change in the DP degree size, the token buffers must be either split or merged accordingly to ensure that the resumed dataloaders do not discard cached data and do not retrain data that has already been sampled and fed. Since sharded states are stored in individual files, ByteCheckpoint can precisely identify the unique items that require resharding and process them efficiently.
@@ 4 Performance Optimization
We now elaborate on ByteCheckpoint’s performance optimization techniques, focusing on minimizing the overhead associated with checkpoint saving, loading, and resharding.
@@ 4.1 Optimized Plan Generation
Plan and metadata cache. In large-scale training, the execution of the planning procedure can introduce significant overhead, particularly when checkpoint saving occurs frequently. For example, planning the saving procedure for a 405B transformer model distributed across 8960 GPUs requires 62 seconds. However, we have observed that both the save plans and the global metadata file remain constant throughout a single training session. This observation presents an opportunity for optimization, as the invariance of these elements allows for potential reuse and caching strategies to reduce overhead in subsequent checkpoint operations. We introduce plan metadata caching, transforming the planning into a one-time cost Once established in the first checkpoint save, the save plans and global metadata file are cached for future reuse, eliminating repetitive planning with communication overhead.
@@ 4.2 Fully Asynchronous Engine Pipeline
ByteCheckpoint Engine optimizes operation execution during checkpoint saving and loading (resharding) through pipelining. Using loading as an example (Fig. 10), we pipeline file reading, deserialization, Host-to-Device (H2D) copy, and inter-GPU communication for each tensor shard, thereby achieving higher efficiency. The read operation downloads files containing desired tensors from the storage system and places them into shared memory (e.g., the dev/shm directory). The deserialize operation deserializes tensors from shared memory. ByteCheckpoint employs multiple threads for parallel file downloading and deserialization. The H2D copy operation transfers tensors from CPU memory to GPUs, while All2All facilitates tensor transferring within the DP group. For saving, we implement a symmetrical, fully asynchronous pipeline comprising Device-to-Host (D2H) copy, serialization, and file uploading operations. To mitigate the performance impact of D2H copy on training, we employ a pinned CPU memory pool combined with a Ping-Pong buffering mechanism to accelerate this operation. We run multiple processes in parallel to serialize tensors and dump files in shared memory. Concurrently, uploading threads actively monitor and initiate file uploading upon the completion of the dumping phase.
@@ 4.3 High Performance Read/Write
We optimize read and write throughput of checkpointing by tailoring our implementation to optimal I/O usage of commonly used storage backends. For example, while HDFS is not primarily designed for random data access, it does offer some random read capabilities via its SDK, allowing applications to access specific file offsets and retrieve data from those positions. We leverage this feature and enable multi-threaded reading of a single file, significantly accelerating the download speed of checkpoint files from HDFS. In our production platform, single-file read speed has improved from 400 MB/s to 2-3 GB/s, all tested on GPU servers equipped with over 100 CPU cores, TB-level memory, and a 200Gbps NIC.
@@
4.4 Decomposing Irregular Tensors
Figure 11: ShardMeta of irregular tensors in optimizer states.
We define irregular tensors as those which, when flattened before sharding, cannot be reshaped back to their original dimensions. These tensors typically arise when applying the Zero Redundancy Optimizer (ZeRO)[33], as implemented in Megatron-LM ZeRO2 and FSDP ZeRO3[57]. In these implementations, optimizer states within a DP group are flattened, concatenated, and sharded. Consequently, the resulting 1D tensor slices often cannot be directly represented using n-dimensional shapes and offsets. Fig. 11 illustrates an example of irregular tensor sharding: Unlike tensors $A$ and $C$, tensor $B$, with an original shape of (3, 2), is evenly split into two shards. Each shard, containing three elements, cannot be directly represented as a two-dimensional shape with corresponding offsets. One intuitive approach to addressing the challenge of irregular tensor shards is to merge all tensor shards into complete tensors before saving checkpoints, thereby simplifying the generation of ShardMeta. For example, to eliminate potential irregular tensors in DCP [46], FSDP performs synchronous all-gather communication operations, interleaved with D2H copy operations for each tensor shard, regardless of whether the shard is irregularly sharded. However, this approach incurs significant communication overhead and requires frequent synchronization between GPU and CPU, which substantially impedes training performance. To mitigate the overhead associated with merging tensor shards, ByteCheckpoint employs a tensor decomposition strategy for managing irregular tensor shards. Specifically, ByteCheckpoint decomposes an irregular tensor into a series of regular ones, representing each with index tuples. For example, consider the irregular tensor shard of tensor $B$ with three elements on rank 00 in Fig. 11. ByteCheckpoint decomposes this shard into two regular tensors that can be directly represented by $nD\_offsets$ and $nD\_lengths$. This approach allows ByteCheckpoint to use multiple ShardMeta entries to represent a single irregular tensor shard. Although such decomposition slightly increases the metadata size and adds steps to the loading procedure, as reconstructing the target tensor may require looking up multiple smaller segments of irregular shards, it avoids the costly communication of tensor shards without extra blocking time during saving.
@@ 5.1 Storage Support
High throughput, scalable storage system. HDFS is the primary storage backend for ByteCheckpoint, we have implemented several optimizations to enhance throughput and scalability for large-scale checkpointing. We have rewritten all HDFS components, including NameNode, DataNode, and SDK [13], in C++, effectively doubling performance compared to the original Java-based implementation. Additionally, we have refined the architecture of HDFS by introducing a new component named NNProxy [1]. NNProxy functions as a stateless RPC proxy for HDFS NameNode, facilitating large-scale federated deployments of NameNodes while maintaining minimal querying latency. The NNProxy design addresses the Queries-Per-Second (QPS) bottleneck of HDFS NameNode metadata requests, a challenge intensified by the high volume of distributed checkpoints. NNProxy also provides additional features such as authentication, rate limiting, and metadata query caching. These capabilities enable more granular management of checkpoints in production environments. Our optimized HDFS achieves a substantial capacity of hundreds of PBs, 10 TB/s read/write bandwidth, and approximately 100,000 metadata QPS.
Checkpoint cool-down strategy. We have implemented a two-tier hot-cold storage architecture using a combination of SSD and HDD storage servers. Our key observation is that newly stored checkpoint files are typically downloaded by evaluation tasks shortly after their creation. However, under normal circumstances where no training anomalies such as loss spikes occur, the access frequency of these files decreases significantly after being downloaded by evaluation tasks. Nonetheless, all checkpoint files must be preserved for traceability. To efficiently manage storage space while ensuring data availability, we have developed a data cool-down mechanism that migrates data from SSD to HDD storage, thus ensuring that high-performance hot storage consistently has sufficient space for operations. Specifically, we cool-down all files that exceed the retention threshold based on their last modification time. We then remap the original file paths to the new HDD storage locations through pure metadata operations. This strategy applies to HDFS directories and preserves the original access paths of the cooled-down files, providing a seamless user experience.

### v3

@@ 3 System Design
Based on the above observations, we design ByteCheckpoint according to the following key principles:
(1) Decoupling. The checkpoint representation remains independent of specific runtime parallelism. Interfaces of training frameworks and storage backends are separated from the core execution engine, ensuring robust extensibility.
(2) User-friendliness. The APIs should be concise, making it seamless for AI researchers and engineers to integrate them into their code and runtime environments.
@@ 3.1 Unified Architecture
The architecture of ByteCheckpoint is illustrated in Fig. 4. Each component is introduced in detail as follows:
API. ByteCheckpoint’s APIs (bytecheckpoint.save and bytecheckpoint.load) provide a unified entrypoint for user code across various training frameworks. For instance, to save checkpoints, users first prepare the corresponding training states, checkpoint path, framework name, and performance-related options, then call bytecheckpoint.save. This high-level entrypoint abstracts underlying system complexities, such as sharding specification, save/reshard plan generation, and I/O operations.
Planner. The Planner serves as the interface for training frameworks. It receives arguments (training states, checkpoint path, etc.) from the API layer, creates ShardMeta (Sec. 3.2) for each tensor shard based on the worker’s rank and framework-specific sharding specification such as Megatron ShardedTensor or FSDP DTensor, and determines the tensors and other states to save/load for each worker. Each worker leverages the Planner to initially create local plans and subsequently collaborates to create global plans. We implement a tailored planner for each training framework to extract information from these specifications and generate plans.
@@
Storage I/O. Analogous to the design of the Planner layer, the Storage I/O layer encapsulates different storage backends and manages backend-specific read/write operations and optimizations. The interface between the Engine layer and the Storage I/O layer remains unified across different storage backends, facilitating the seamless integration of new storage backends. ByteCheckpoint supports multiple storage options, including in-memory checkpoint storage [66], local disk storage, and remote storage systems.
@@
Fig. 5 illustrates a typical use case of ByteCheckpoint. Initially, users define a dictionary specifying the states to be saved or loaded, encompassing the model, optimizer, dataloader, and additional states. At the outset of training resumption, stage transition, or evaluation, bytecheckpoint.load is invoked to load saved checkpoints. Checkpoint resharding occurs automatically during loading when there are changes in parallelism. During training, bytecheckpoint.save is periodically invoked to save checkpoints.
@@ 3.2 Decoupled Checkpoint Representation
Model and optimizer states. Each tensor is uniquely identified by a "fully qualified name" (FQN) and has a global shape, representing its original shape before sharding. Tensors can be either sharded or replicated across ranks (aka training workers). For sharded tensors, the specific shard held by a worker is determined by three factors: the parallelism (e.g., tensor parallelism) applied to the tensor, the sharding dimension, and the group rank within the corresponding parallelism group. Based on a tensor’s FQN, global shape, and sharding specification, ByteCheckpoint creates the tensor shard metadata (ShardMeta) for each rank, which is utilized in checkpoint storage to represent the tensor shard, independent of parallelism. More precisely, ShardMeta for a tensor shard is an index tuple (fqn, nD_offsets, nD_lengths), where nD_offsets and nD_lengths indicate the local shard’s offsets and lengths along the original multi-dimensional axes in its global shape.
Figure 6: Checkpoint representation in ByteCheckpoint. The pipeline parallelism is applied in this example.
@@
Checkpoint representation. Fig. 6 illustrates the checkpoint representation in ByteCheckpoint. Distributed checkpoints comprise a global metadata file and multiple storage files. Each rank generates three distinct files: a model state file, an optimizer state file, and an extra state file. The Dataloader state file is generated only by training workers whose ranks for all parallelism degrees, except for DP degrees, are 0. For tensor shards in model and optimizer states, their Metadata consists of three parts: BasicMeta, which records essential information of individual tensor shards such as stride and device, critical for recovering the runtime state; ShardMeta, as previously introduced, recording relative position information of shards in the complete tensor; ByteMeta, which specifies the byte start offset and length of each tensor shard within the storage file. All tensor metadata are consolidated into the global metadata file, and a mapping, termed TensorShardtoBasicByteMap, is established between saved tensor shards and storage files based on the ShardMeta, BasicMeta, and ByteMeta of each tensor shard, ensuring accurate data retrieval. The global metadata file also includes a LoaderShardtoByteMap, which records the file index information of sharded states in each dataloader. All storage files and the global metadata file are stored in the specified storage backend designated by the checkpoint path.
Figure 7: ShardMeta of irregular tensors in optimizer states.
Decomposing irregular tensors. We define irregular tensors as those which, when flattened before sharding, cannot be reshaped back to their original dimensions. These tensors typically arise when applying the Zero Redundancy Optimizer (ZeRO)[42], as implemented in Megatron-LM ZeRO2 and FSDP ZeRO3[73]. In these implementations, optimizer states within a DP group are flattened, concatenated, and sharded. Consequently, the resulting 1D tensor slices often cannot be directly represented using n-dimensional shapes and offsets. Fig. 7 illustrates an example of irregular tensor sharding: Unlike tensors $A$ and $C$, tensor $B$, with an original shape of (3, 2), is evenly split into two shards. Each shard, containing three elements, cannot be directly represented as a two-dimensional shape with corresponding offsets. One intuitive approach to addressing the challenge of irregular tensor shards is to merge all tensor shards into complete tensors before saving checkpoints, thereby simplifying the generation of ShardMeta. For example, to eliminate potential irregular tensors in DCP [58], FSDP performs synchronous all-gather communication operations, interleaved with D2H copy operations for each tensor shard, regardless of whether the shard is irregularly sharded. However, this approach incurs significant communication overhead and requires frequent synchronization between GPU and CPU, substantially impeding efficiency.
To mitigate the overhead associated with merging tensor shards, ByteCheckpoint employs a tensor decomposition strategy for managing irregular tensor shards. Specifically, ByteCheckpoint decomposes an irregular tensor into a series of regular ones, representing each with index tuples. For example, consider the irregular tensor shard of tensor $B$ with three elements on rank 00 in Fig. 7. ByteCheckpoint decomposes this shard into two regular tensors that can be directly represented by $nD\_offsets$ and $nD\_lengths$. This approach allows ByteCheckpoint to use multiple ShardMeta entries to represent a single irregular tensor shard. Although such decomposition slightly increases the metadata size and adds steps to the loading procedure, as reconstructing the target tensor may require looking up multiple smaller segments of irregular shards, it avoids the costly communication of tensor shards without extra blocking time during saving.
@@ 3.3 Checkpoint Resharding Workflow
Figure 8: Load-time tensor resharding workflow. Checkpoint $k$ denotes the checkpoint saved by a previous live worker $k$.
ByteCheckpoint implements a generic workflow for saving and loading, automatically resharding checkpoints during the loading process. Leveraging the independence between the Engine Layer, Planner Layer, and Storage I/O Layer, we can execute consistent saving and loading steps across different training frameworks and storage backends. We take tensor resharding as an example in Fig. 8, emphasizing how our checkpoint representation enables flexible load-time resharding. The process for checkpoint saving and loading without resharding follows a similar procedure.
Step 1. To initiate load-time checkpoint resharding, each rank invokes the bytecheckpoint.load() API, specifying the checkpoint path along with the model/optimizer to be restored. All ranks then load the global metadata file from the path.
Step 2. For each tensor shard in the given model/optimizer, each rank queries the TensorShardToBasicByteMap within the global metadata file, identifying matching segments between the saved tensor shards and the sharding specification of new shards. After identifying these matches, the planner constructs a local load plan, which includes BasicMeta and ByteMeta tailored to the specific shard. This identification mechanism is illustrated at the bottom of Fig. 8.
@@
Step 5. The execution engine on each rank selects the storage backend wrapper in the Storage I/O layer based on the checkpoint path and then executes the loading pipeline (Sec. 4.2).
Step 6. Upon completion of the loading pipeline, each rank utilizes the optimized asynchronous collective barrier primitive ensuring the atomicity of the distributed loading. Further details can be found in Appendix B.
Dataloader resharding. Similar to tensor resharding in Fig 8, ByteCheckpoint reshards the dataloader by querying the LoaderShardtoByteMap. The illustration of dataloader resharding is depicted in Fig. 9. When parallelism configuration changes, for the dataloader checkpoints, common items in the dataloader checkpoints can be loaded directly, whereas unique items such as accumulative token buffers and data retrieval offsets require resharding. Specifically, when the DP degree size remains constant while other parallel degrees are altered (TP degree changes in the example of Fig. 9), the token buffers should be copied to the destination workers for bitwise-correct resuming. when there is a change in the DP degree size, the token buffers must be either split or merged accordingly to ensure that the resumed dataloaders do not discard cached data and do not retrain data that has already been sampled and fed. Thanks to the split storage strategy designed for the dataloader module, ByteCheckpoint can precisely identify the unique items that require resharding and process them efficiently.
Figure 9: Examples of dataloader resharding. Sharded states such as cumulative token buffers are resharded according to the changes of parallelism configurations. We do not depict data retrieval offsets for clarity of the figure.
@@ 4 Performance Optimization
We now elaborate on ByteCheckpoint’s performance optimization techniques, focusing on minimizing the overhead associated with checkpoint saving and loading.
@@ 4.1 Optimized Plan Generation
Plan and metadata cache. In large-scale training, the execution of the planning procedure can introduce significant communication overhead, particularly when checkpoint saving occurs frequently. For example, planning the saving procedure for a 405B transformer model distributed across 8960 GPUs requires 62 seconds. However, we have observed that both the save plans and the global metadata file, although coupled with specific parallelism, remain constant throughout a single training session. This allows for caching strategies to reduce planning times and associated overheads in subsequent checkpoint operations. We introduce plan metadata caching, transforming the planning into a one-time cost. Once established for the first time, the save plans and global metadata file are cached for future reuse, eliminating repetitive planning.
@@ 4.2 Fully Asynchronous Engine Pipeline
ByteCheckpoint Engine optimizes operation execution during checkpoint saving and loading (resharding) through pipelining. Using loading as an example (Fig. 10), we pipeline file reading, deserialization, Host-to-Device (H2D) copy, and inter-GPU communication for each tensor shard, thereby achieving higher efficiency. The read operation downloads files containing desired tensors from the storage system and places them into shared memory (e.g., the dev/shm directory). The deserialize operation deserializes tensors from shared memory. ByteCheckpoint employs multiple threads for parallel file downloading and deserialization. The H2D copy operation transfers tensors from CPU memory to GPUs, while All2All facilitates tensor transferring within the DP group. For saving, we implement a symmetrical, fully asynchronous pipeline comprising D2H copy, serialization, and file uploading operations. To mitigate the performance impact of D2H copy on training, we employ a pinned CPU memory pool combined with a Ping-Pong buffering mechanism to accelerate this operation. We run multiple parallel processes to serialize tensors and dump files into shared memory. Uploading threads actively monitor and initiate file uploading upon the completion of the dumping phase.
@@ 4.3 High Performance Read/Write
We optimize the read and write throughput of checkpointing by tailoring our implementation to optimal I/O usage of commonly used storage backends. For example, while HDFS is not primarily designed for random data access, it does offer some random read capabilities via its SDK, allowing applications to access specific file offsets and retrieve data from those positions. We leverage this feature and enable multi-threaded reading of a single file, significantly accelerating the download speed of checkpoint files from HDFS. In our production platform, single-file read speed has improved from 400 MB/s to 2-3 GB/s, all tested on H800 servers equipped with over 100 CPU cores, TB-level memory, and a 200Gbps NIC.
@@ 5.1 Storage Support
High throughput, scalable storage system. HDFS is the primary storage backend for ByteCheckpoint, we have implemented several optimizations to enhance throughput and scalability for large-scale checkpointing. We have rewritten all HDFS components, including NameNode, DataNode, and SDK [15], in C++, effectively doubling performance compared to the original Java-based implementation. Additionally, we have refined its architecture by introducing a new component named NNProxy [1]. NNProxy functions as a stateless RPC proxy for HDFS NameNode, facilitating large-scale federated deployments of NameNodes while maintaining minimal querying latency. The NNProxy design addresses the QPS bottleneck of HDFS NameNode metadata requests, a challenge intensified by the high volume of distributed checkpoints. NNProxy also provides additional features such as authentication, rate limiting, and metadata query caching. These capabilities enable more granular management of checkpoints in production environments. Our optimized HDFS achieves a substantial capacity of hundreds of PBs, 10 TB/s read/write bandwidth, and approximately 100,000 metadata QPS.
Checkpoint cool-down strategy. We have implemented a two-tier hot-cold storage architecture using a combination of SSD and HDD storage servers. Our key observation is that newly stored checkpoint files are typically downloaded by evaluation tasks shortly after their creation. However, where no training anomalies such as loss spikes occur, the access frequency of these files decreases significantly after being downloaded by evaluation tasks. Nonetheless, all checkpoint files must be preserved for traceability. To efficiently manage storage space while ensuring data availability, we have developed a data cool-down mechanism that migrates data from SSD to HDD storage, thus ensuring that high-performance hot storage consistently has sufficient space for operations. Specifically, we cool down all files that exceed the retention threshold based on their last modification time. We then remap the original file paths to the new HDD storage locations through pure metadata operations. This strategy applies to HDFS directories and preserves the original access paths of the cooled-down files, providing a seamless user experience.